# Is Your HD Map Constructor Reliable under Sensor Corruptions?

### v1

@@ 1 Introduction
HD maps are fundamental components in autonomous driving systems, providing centimeter-level details of traffic rules, vectorized topology, and navigation information [31; 38]. These maps enable the ego-vehicle to accurately locate itself on the road and anticipate upcoming features [36; 69; 11; 45]. HD map constructors formulate this task as predicting a collection of vectorized static map elements in bird’s eye view (BEV), e.g., pedestrian crossings, lane dividers, road boundaries, etc. [37; 68; 72].
Existing HD map construction methods can be categorized based on the input sensor modality: camera-only [31; 38; 36; 69; 45] LiDAR-only [31; 38; 36; 37], and camera-LiDAR fusion [31; 38; 36] models. Each sensor poses distinct functionalities: cameras capture semantic-rich information from images, while LiDAR provides explicit geometric information from point clouds [37; 43; 23]. Generally, camera-based methods outperform LiDAR-only methods, and fusion-based methods yield the most satisfactory results [27; 72]. However, current model designs and performance evaluations are based on ideal driving conditions, e.g., clear daytime weather and fully functional sensors [3; 30]. In real-world driving scenarios, adverse conditions, such as bad weather, motion distortions, and sensor malfunctions (frame loss, sensor crashes, incomplete echoes, etc.) are unavoidable [57; 24]. It remains unclear how existing HD map construction methods perform under such challenging yet safety-critical conditions, highlighting the need for a thorough out-of-domain robustness evaluation.
@@
1) Among all camera/LiDAR corruptions, Snow corruption significantly degrades model performance; it covers the road, rendering map elements unrecognizable and posing a major threat to autonomous driving. Besides, sensor failure corruptions (e.g.. Frame Lost and Incomplete Echo) are also challenging for all models, demonstrating the serious threats of sensor failures on HD map models.
2) While Camera-LiDAR fusion methods have shown promising performance by incorporating information from both modalities [36; 72], existing methods often assume access to complete sensor information, leading to poor robustness and potential collapse when sensors are corrupted or missing.
@@ 2 Related Work
HD Map Construction. The construction of HD maps is a critical yet extensively researched area. Based on the input sensor modality, existing literature can be categorized into camera-only [36; 69; 11; 45; 37; 68; 72], LiDAR-only [31; 38], and camera-LiDAR fusion [36; 37; 72] models. Camera-based methods [31; 38; 36; 69; 11; 45; 37; 68] have increasingly employed the BEV representation as an ideal feature space for multi-view perception due to its ability to mitigate scale ambiguity and occlusion challenges. Techniques such as LSS [44], Deformable Attention [34], and GKT [5] have been proposed to project perspective view (PV) features into the BEV space by leveraging geometric priors. However, these methods lack explicit depth information. Consequently, they have come to rely on higher resolution images or larger backbones to achieve enhanced accuracy [40; 39; 53; 34; 64; 60]. LiDAR-based methods [51; 31; 38; 37; 36] benefit from the accurate 3D geometric information provided by LiDAR inputs but face challenges related to data sparsity and sensing noise.
@@
Multi-Sensor HD Map Construction. Camera-LiDAR fusion-based methods can be roughly divided into three categories: point-level fusion [49; 48; 61; 8; 65], feature-level fusion [66; 1; 6; 7; 35], and BEV-level fusion [36; 37; 72]. Recently, camera-LiDAR feature fusion in the unified BEV space has gained attention. BEV-level fusion integrates features from camera and LiDAR sensors within the same BEV space, combining complementary modalities to achieve superior performance over uni-modal approaches. Despite the progress in HD map construction methods, their resilience to real-world challenges such as adverse weather and sensor failures remains unclear, raising safety concerns [25; 26]. In this work, we make the first attempt to explore the robustness of existing HD map construction methods under sensor corruptions that occur in real-world environments.
@@
Robustness against Sensor Corruptions. Assessing the robustness of driving perception models under sensor corruptions has emerged as a crucial research area [12; 73; 25; 57; 26]. Recently, the corruption robustness of BEV perception tasks has been extensively studied. RoboDepth [25; 46] establishes a robustness benchmark for monocular depth estimation under corruptions, while RoboBEV [57; 58] introduces a comprehensive benchmark for evaluating the robustness of four BEV perception tasks, including 3D object detection [34; 41], semantic segmentation [70; 71], depth estimation [54], and semantic occupancy prediction [55; 21]. However, RoboBEV’s analyses of multi-modal fusion model robustness only consider complete sensor failure, overlooking other common sensor corruptions and their combinations. Dong et al. [12] systematically design $27$ types of common corruptions for 3D object detection in both LiDAR and camera sensors. Meanwhile, Robo3D [24] benchmarks the robustness of 3D detectors and segmentors against LiDAR corruptions.
@@
Comparison with Existing Works. This work differs from prior literature in three key aspects. Firstly, we focus on the vectorized HD map construction task, distinct from other BEV perception tasks [12; 25; 57]. Secondly, we introduce new sensor corruption types that closely mimic real-world scenarios. Specifically, we design $13$ new multi-sensor corruption types to benchmark camera-LiDAR fusion models comprehensively, surpassing the scope of complete sensor failure analysis in RoboBEV [57]. Thirdly, we explore distinct data augmentation techniques that are applied to LiDAR point clouds and RGB images to analyze their impact on enhancing corruption robustness. To the best of our knowledge, MapBench serves as the first study to comprehensively benchmark and evaluate the reliability of HD map construction methods again single- and multi-modal sensor corruptions.
@@ 3.2 Evaluation Metrics
Inspired by  [17; 24; 57], we define two robustness evaluation metrics based on mAP (mean Average Precision), a commonly-used accuracy indicator for vectorized HD map construction.
@@ 4.1 Benchmark Configuration
Candidate Models. Our MapBench encompasses a total of $31$ HD map constructors and their variants, i.e., HDMapNet [31], VectorMapNet [38], PivotNet [11], BeMapNet [45], MapTR [36], MapTRv2 [37], StreamMapNet [68] and HIMap [72]. The code of some other HD map methods [32; 59; 19; 69; 62; 42; 4; 22; 60; 67; 50] are not open source, thus will not be considered in this work.
@@
((a)) Corrupt-C + Clean-L
((b)) Corrupt-L + Clean-C
((c)) Corrupt-C + Corrupt-L
Figure 4: The results of Camera-LiDAR fusion methods [36; 72] under multi-sensor corruptions.
@@ 5 Observation & Discussion
Backbones. We first comprehensively investigate the impact of backbone networks, with results presented in Tab. 5. Specifically, we use three different backbones in PivotNet [11] and BeMapNet [45], respectively. The results demonstrate that Swin Transformer [40] significantly retains model robustness. As an example, compared with ResNet-50 [15], the Swin Transformer [40] backbone improves the mCE of PivotNet [11] and BeMapNet [45] with $22.2\%$ and $24.1\%$ absolute gains, respectively. The results demonstrate that larger pretrained models tend to help enhance the robustness of feature extraction under out-of-domain data, which is in line with the observation drawn in [16; 2; 13; 10; 52].

### v3

@@ 1 Introduction
HD maps are fundamental components in autonomous driving systems, providing centimeter-level details of traffic rules, vectorized topology, and navigation information [35, 43]. These maps enable the ego-vehicle to accurately locate itself on the road and anticipate upcoming features [41, 74, 11, 50]. HD map constructors formulate this task as predicting a collection of vectorized static map elements in bird’s eye view (BEV), e.g., pedestrian crossings, lane dividers, road boundaries, etc. [42, 73, 77].
Existing HD map construction methods can be categorized based on the input sensor modality: camera-only [35, 43, 41, 74, 50] LiDAR-only [35, 43, 41, 42], and camera-LiDAR fusion [35, 43, 41] models. Each sensor poses distinct functionalities: cameras capture semantic-rich information from images, while LiDAR provides explicit geometric information from point clouds [42, 48, 26, 29]. Generally, camera-based methods outperform LiDAR-only methods, and fusion-based methods yield the most satisfactory results [31, 77]. However, current model designs and performance evaluations are based on ideal driving conditions, e.g., clear daytime weather and fully functional sensors [3, 34]. In real-world driving scenarios, adverse conditions, such as bad weather, motion distortions, and sensor malfunctions (frame loss, sensor crashes, incomplete echoes, etc.) are unavoidable [62, 27]. It remains unclear how existing HD map construction methods perform under such challenging yet safety-critical conditions, highlighting the need for a thorough out-of-domain robustness evaluation.
@@
1) Among all camera/LiDAR corruptions, Snow corruption significantly degrades model performance; it covers the road, rendering map elements unrecognizable and posing a major threat to autonomous driving. Besides, sensor failure corruptions (e.g., Frame Lost and Incomplete Echo) are also challenging for all models, demonstrating the serious threats of sensor failures on HD map models.
2) While Camera-LiDAR fusion methods have shown promising performance by incorporating information from both modalities [41, 77], existing methods often assume access to complete sensor information, leading to poor robustness and potential collapse when sensors are corrupted or missing.
@@ 2 Related Work
HD Map Construction. The construction of HD maps is a critical yet extensively researched area. Based on the input sensor modality, existing literature can be categorized into camera-only [41, 74, 11, 50, 42, 73, 77, 15], LiDAR-only [35, 43], and camera-LiDAR fusion [41, 42, 77] models. Camera-based methods [35, 43, 41, 74, 11, 50, 42, 73] have increasingly employed the BEV representation as an ideal feature space for multi-view perception due to its ability to mitigate scale ambiguity and occlusion challenges. Techniques such as LSS [49], Deformable Attention [39], and GKT [5] have been proposed to project perspective view (PV) features into the BEV space by leveraging geometric priors. However, these methods lack explicit depth information. Consequently, they have come to rely on higher resolution images or larger backbones to achieve enhanced accuracy [45, 44, 58, 39, 69, 65]. LiDAR-based methods [56, 35, 43, 42, 41, 38] benefit from the accurate 3D geometric information provided by LiDAR inputs but face challenges related to data sparsity and sensing noise.
@@
Multi-Sensor HD Map Construction. Camera-LiDAR fusion-based methods can be roughly divided into three categories: point-level fusion [54, 53, 66, 8, 70], feature-level fusion [71, 1, 6, 7, 40], and BEV-level fusion [41, 42, 77, 17]. Recently, camera-LiDAR feature fusion in the unified BEV space has gained attention. BEV-level fusion integrates features from camera and LiDAR sensors within the same BEV space, combining complementary modalities to achieve superior performance over uni-modal approaches. Despite the progress in HD map construction methods, their resilience to real-world challenges such as adverse weather and sensor failures remains unclear, raising safety concerns [28, 30]. In this work, we make the first attempt to explore the robustness of existing HD map construction methods under sensor corruptions that occur in real-world environments.
@@
Robustness against Sensor Corruptions. Assessing the robustness of driving perception models under sensor corruptions has emerged as a crucial research area [16, 12, 78, 28, 62, 30]. Recently, the corruption robustness of BEV perception tasks has been extensively studied. RoboDepth [28, 51] establishes a robustness benchmark for monocular depth estimation under corruptions, while RoboBEV [62, 63] introduces a comprehensive benchmark for evaluating the robustness of four BEV perception tasks, including 3D object detection [39, 46], semantic segmentation [75, 76], depth estimation [59], and semantic occupancy prediction [60, 24]. However, RoboBEV’s analyses of multi-modal fusion model robustness only consider complete sensor failure, overlooking other common sensor corruptions and their combinations. Dong et al. [12] systematically design $27$ types of common corruptions for 3D object detection in both LiDAR and camera sensors. Meanwhile, Robo3D [27] benchmarks the robustness of 3D detectors and segmentors against LiDAR corruptions.
@@
Comparison with Existing Works. This work differs from prior literature in three key aspects. Firstly, we focus on the vectorized HD map construction task, distinct from other BEV perception tasks [12, 28, 62]. Secondly, we introduce new sensor corruption types that closely mimic real-world scenarios. Specifically, we design $13$ new multi-sensor corruption types to benchmark camera-LiDAR fusion models comprehensively, surpassing the scope of complete sensor failure analysis in RoboBEV [62]. Thirdly, we explore distinct data augmentation techniques that are applied to LiDAR point clouds and RGB images to analyze their impact on enhancing corruption robustness. To the best of our knowledge, MapBench serves as the first study to comprehensively benchmark and evaluate the reliability of HD map construction methods again single- and multi-modal sensor corruptions.
@@ 3.2 Evaluation Metrics
Inspired by  [20, 27, 62], we define two robustness evaluation metrics based on mAP (mean Average Precision), a commonly-used accuracy indicator for vectorized HD map construction.
@@ 4.1 Benchmark Configuration
Candidate Models. Our MapBench encompasses a total of $31$ HD map constructors and their variants, i.e., HDMapNet [35], VectorMapNet [43], PivotNet [11], BeMapNet [50], MapTR [41], MapTRv2 [42], StreamMapNet [73] and HIMap [77]. The code of some other HD map methods [36, 64, 22, 74, 67, 47, 4, 25, 65, 72, 55] are not open source, thus will not be considered in this work.
@@
(a) Corrupt-C + Clean-L
(b) Corrupt-L + Clean-C
(c) Corrupt-C + Corrupt-L
Figure 4: The results of Camera-LiDAR fusion methods [41, 77] under multi-sensor corruptions.
@@ 5 Observation & Discussion
Backbones. We first comprehensively investigate the impact of backbone networks, with results presented in Tab. 5. Specifically, we use three different backbones in PivotNet [11] and BeMapNet [50], respectively. The results demonstrate that Swin Transformer [45] significantly retains model robustness. As an example, compared with ResNet-50 [18], the Swin Transformer [45] backbone improves the mCE of PivotNet [11] and BeMapNet [50] with $22.2\%$ and $24.1\%$ absolute gains, respectively. The results demonstrate that larger pretrained models tend to help enhance the robustness of feature extraction under out-of-domain data, which is in line with the observation drawn in [19, 2, 13, 10, 57].